OpenStudy (noobster916):

how many grams are in 2.3 x 10^-4 moles of calcium phosphate, using scientific terms, Ca^3(PO^3)^2

OpenStudy (noobster916):

My teacher is asian and we cant understand a word she's saying so im lost 100%

OpenStudy (noobster916):

Ca^3 (Ca x 3) (P (P x 1) O^3 (O x 3)) x 2

OpenStudy (noobster916):

Is the mass for each element shown

OpenStudy (dallascowboys88):

Yes Your Right

OpenStudy (dallascowboys88):

Yes They Are Shown.

OpenStudy (noobster916):

...

OpenStudy (noobster916):

So, Ca ^ 3 (Calcium = 40 x 3 = 120) 120 grams so far! Then, it gets to the difficult part. So, PO^3 is basically saying P (Phosphorus = 31) So 31 so far ( Don't add up ) And O ( Oxygen = 16 x 3 = 48) Then you can do 31 + 48 = 79, which means you have to multiply 79 by two I believe, meaning it will be 158 So, 158 + 120 = 278 However, this could be incorrect as it may be 79 squared, so 79 x 79 = 6241 + 120 = 6361, and if we look at 2.3 x 10 ^ -4, In scientific notation, it basically means 0.0023 I believe... So 6361 and 0.0023 (-4) How many grams?